# Newcastle-upon-Tyne, UK

### v3

@@ 4.3 Multi-modal Sensors
Many sensor-based security proposals [13, 12, 4, 2, 3] assert that combining multiple sensor modalities can bolster security, based on the intuition that an adversary must accurately predict several data streams, rather than just one. A naïve model assumes each sensor is statistically independent, yielding the simplified relation $H(X_{1},X_{2},\dots,X_{n})=\sum_{i=1}^{n}H(X_{i})$. However, real-world sensor outputs often exhibit significant interdependence. For example, linear acceleration, rotation vector, and gravity in the Android ecosystem are often derived from the same underlying motion sensors, causing strong correlations. Figure 2 illustrates the correlation matrices of the sensors we examine across our four datasets, revealing that many modalities are far from independent.
@@
Unfortunately, directly estimating the joint entropy of discrete sensor data often becomes computationally infeasible as the number of modalities increases, due to the curse of dimensionality. Specifically, for $n$ modalities, where each modality $X_{i}$ is discretised into $b_{i}$ bins (with $\{b_{1},b_{2},\dots,b_{n}\}\subseteq B$), the total number of joint bins scales as $\prod^{n}_{i=1}b_{i}$. Applying Freedman–Diaconis binning rules typically results in thousands of bins per modality, causing the number of joint bins to explode combinatorially. Consequently, computing a full joint entropy estimation, as in Eq. 6, becomes intractable. Another concern is the combinatorial explosion of sensor combinations. Let us assume that one has 19 modalities, as in Table 1, then the total number of combinations to analyse is equal to $202^{19}-$ 20, skipping single sensors and the empty set. For even moderately large datasets, iterating over every sensor subset and computing its full joint entropy is prohibitively expensive. Addressing these problems is challenging. One mitigation is to reduce the number of bins per modality. However, this inevitably risks oversimplifies the distribution and underestimating its joint entropy. Preliminary experiments confirmed joint entropies could be computed directly for $n\leq 3$ modalities with a maximum of 1250 bins and fewer than 150K total samples from the Relay dataset. Further scaling became infeasible. We also observed that imposing these smaller bin counts (rather than the thousands implied by Freedman–Diaconis) reduced our single-sensor entropy estimates by approximately 2–3 bits on average compared to those reported in Table 1.
@@ 5.1 Comparison with Related Work
Multiple sensors improve worst-case entropy, but not dramatically. While combining more modalities typically raises the upper bound (e.g. Hartley or Shannon entropy), our experiments reveal that min-entropy gains are far smaller than one might hope. In some cases, $H_{\infty}$ plateaus, indicating that a few highly probable outcomes still dominate the distribution.
Overall, our findings place stricter bounds on the real benefits of mobile-sensor-based approaches. While sensor data can be effective for classification-based tasks such as continuous authentication, where strong assurances are not necessarily required, or simple proximity checks, the suitability of sensor data as an attack-resistant random source remains doubtful unless carefully engineered and rigorously evaluated. Min-entropy is critical in scenarios where an adversary can guess the most likely states first; our analysis does not even incorporate adversarial perturbations, meaning real-world attacks could degrade unpredictability even further. Hence, while mobile sensors can augment other authentication or key-generation processes, they rarely suffice as a standalone replacement for dedicated, cryptographically validated entropy sources.
@@ 5.1.1 Can Extractors Salvage Weak Sources?
One might hope that cryptographic extractors (e.g. Von Neumann extractors or more advanced schemes [46, 47, 48]) could mitigate low-entropy sensor data. Indeed, extractors are designed to reduce bias in a noisy or skewed source. However, they cannot increase the total amount of randomness beyond the source’s intrinsic min-entropy. In other words, if the combined sensor data only provides around 24 bits of min-entropy, then post-processing can at best produce a short, unbiased bitstring reflecting those 24 bits, and no more. From a security standpoint, this means that an extractor can improve the quality of the randomness (i.e., make it more uniform) but not increase the quantity (i.e., its total brute-force resistance). Thus, while extraction can help if the original data has enough min-entropy to meet system requirements, it cannot elevate a weak source—one that fundamentally lacks sufficient min-entropy—to the higher security thresholds often demanded by rigorous standards.
@@ 5.2 Limitations
Despite our comprehensive multivariate analysis, this work has some limitations. Firstly, although we analyse four large datasets, they do not fully capture contexts all possible contexts. For example, work has suggested that performing dedicated movements (e.g. gestures) can increase the amount of usable entropy from motion sensors, on the order of magnitude of 5–6 bits [25, 22]. Our datasets do not cover such dedicated movements; it is possible that the reported resuts are an underestimation of entropy for motion-based sensors. Secondly, our choice of Freedman-Diaconis binning and Chow-Liu trees is a pragmatic compromise. Smaller bin sizes tend to over-simplify the distribution and underestimate entropy, whereas larger bins can lead to computational blowup. Although our approach outperforms naive joint entropy estimation, it is still an approximation. Thirdly, we largely focus on a data-centric view of entropy and do not deeply explore sophisticated adversarial strategies, such as sensor spoofing [3], fault injection attacks [34] or cross-modality correlation attacks (e.g. see [49] for such attacks against wireless body area networks). These threats may significantly reduce the usable entropy of sensors even further.
@@ 6 Conclusion
Our findings also challenge the prevailing evaluation approach that machine learning metrics (e.g., accuracy or EER) suffice to demonstrate the inherent randomness of sensor signals. The vulnerability of mobile sensors to biased distributions significantly undermine their effectiveness as reliable entropy sources. We also cast doubt on the use of sensors with respect to their non-stationary and lack of reproducibility. These issues collectively contradict the criteria articulated in frameworks such as NIST SP 800-90B, which emphasise noise-source stationarity and protection from external influence. The effectiveness of countermeasures remains an open research challenge. Our hope is that the methodologies and insights presented here will encourage the security community to adopt more rigorous evaluation strategies for sensor-based techniques, and ultimately pave the way for safer and more robust designs in mobile-device security.

### v4

@@ 4.3 Multi-modal Sensors
Several sensor-based security proposals [13, 12, 4, 2, 3] assert that combining multiple sensor modalities can bolster security, based on the intuition that an adversary must accurately predict several data streams, rather than just one. This section will examine that claim.
A naïve approach to analysing the entropy of multiple sensors is to first assume that each sensor is statistically independent, yielding the simplified relation $H(X_{1},X_{2},\dots,X_{n})=\sum_{i=1}^{n}H(X_{i})$. However, real-world sensor outputs often exhibit significant interdependence. For example, the linear acceleration, rotation vector, and gravity sensors in the Android ecosystem can be implemented in software using data from accelerometers and gyroscopes, creating a source of correlation [44]. Figure 2 shows the correlation matrices of the modalities across our four datasets, revealing that many modalities are far from independent.
@@
Unfortunately, directly estimating the joint entropy of discrete sensor data becomes computationally infeasible as the number of modalities increases due to the curse of dimensionality. Specifically, for $n$ modalities, where each modality $X_{i}$ is discretised into $b_{i}$ bins (with $\{b_{1},b_{2},\dots,b_{n}\}\subseteq B$), the total number of joint bins scales as $\prod^{n}_{i=1}b_{i}$. Applying Freedman–Diaconis binning rules typically results in thousands of bins per modality, causing the number of joint bins to explode combinatorially. Consequently, computing a full joint entropy estimation, as in Eq. 6, becomes intractable. Another concern is the combinatorial explosion of sensor combinations. Let us assume that one has 19 modalities, as in Table 1, then the total number of combinations to analyse is equal to $202^{19}-$ 20, skipping single sensors and the empty set. For even moderately large datasets, iterating over every sensor subset and computing its full joint entropy is prohibitively expensive.
Addressing these problems is challenging. One mitigation is to reduce the number of bins per modality. However, this inevitably risks oversimplifying the distribution and underestimating its joint entropy. Preliminary experiments confirmed joint entropies could be computed directly for $n\leq 3$ modalities with a maximum of 1250 bins and fewer than 150K total samples from the Relay dataset. Further scaling became infeasible. We also observed that imposing these smaller bin counts (rather than the thousands implied by Freedman–Diaconis) reduced our single-sensor entropy estimates by approximately 2–3 bits on average compared to those reported in Table 1.
@@ 5.1 Comparison with Related Work
Multiple sensors improve worst-case entropy, but not dramatically. While combining more modalities typically raises the upper bound (e.g. Hartley or Shannon entropy), our experiments reveal that min-entropy gains are far smaller than one might hope. In some cases, $H_{\infty}$ plateaus, indicating that a few highly probable outcomes still dominate the distribution. Min-entropy is critical in scenarios where an adversary can guess the most likely states first. Overall, our findings place stricter bounds on the benefits and risks of mobile-sensor-based approaches. We estimate that even the most complicated multi-modal combination provide relatively little entropy. It is important to note that our analysis does not even incorporate adversarial perturbations, meaning real-world attacks could degrade unpredictability even further. Hence, while mobile sensors can augment other authentication or key-generation processes, they rarely suffice as a standalone source. They might be useful for tasks where strong assurances are not required, e.g. simple proximity checks; however, the suitability of sensor data as a robust source of unpredictability remains highly doubtful.
@@ 5.1.1 Can Extractors Salvage Weak Sources?
One might hope that cryptographic extractors (e.g. Von Neumann extractors or more advanced schemes [47, 48, 49]) could mitigate low-entropy sensor data. Extractors are designed to reduce bias in a noisy or skewed source; however, they cannot increase the total amount of randomness beyond the source’s intrinsic min-entropy. That is, if the combined sensor data only provides around 24 bits of min-entropy, then post-processing can at best produce a short, unbiased bitstring reflecting those 24 bits, and no more. From a security standpoint, this means that an extractor can improve the quality of the randomness (i.e., make it more uniform) but not increase the quantity (i.e., its total brute-force resistance). Thus, while extraction can help if the original data has enough min-entropy to meet system requirements, it cannot elevate a weak source—one that fundamentally lacks sufficient min-entropy—to the higher security thresholds often demanded by rigorous standards.
@@ 5.2 Limitations
Despite our comprehensive analysis, this work has some limitations. Firstly, while we analyse four large datasets, they do not fully capture contexts all possible contexts. Some work has suggested that performing dedicated movements (e.g. gestures) can increase the amount of usable entropy from motion sensors by 5–6 bits [26, 22]. Our datasets do not cover such dedicated movements; it is possible that the reported results are an underestimation of entropy for sensors which are deliberately perturbed as an entropy-generating action. Secondly, our choice of Freedman-Diaconis binning and Chow-Liu trees is a pragmatic compromise. Smaller bin sizes tend to over-simplify the distribution and underestimate entropy, whereas larger bins can lead to computational blowup. As such, although our approach outperforms naïve joint entropy estimation, it is still an approximation. Thirdly, we largely focus on a data-centric view of entropy and do not examine other threats, such as sensor spoofing [3], hardware attacks [34] and cross-device correlation attacks (see [50] against wireless body area networks). It is possible that these threats may reduce the entropy of sensors even further.
@@ 6 Conclusion
Our findings also challenge the prevailing orthodoxy that model evaluation metrics (e.g. accuracy or EER) suffice to demonstrate the inherent randomness of sensor signals. The vulnerability of mobile sensors to biased distributions significantly undermine their effectiveness as reliable entropy sources. We also cast doubt on the use of sensors with respect to their non-stationary and lack of reproducibility. These issues collectively contradict the criteria articulated in frameworks such as NIST SP 800-90B, which emphasise noise-source stationarity and protection from external influence. The effectiveness of countermeasures remains an open research challenge. Our hope is that the methodologies and insights presented here will encourage the security community to adopt more rigorous evaluation strategies for sensor-based techniques, and ultimately pave the way for safer and more robust designs in mobile-device security.